# Pretraining Scaling Laws for Generative Evaluations of Language Models

### v1

@@ Abstract
Neural scaling laws have played a central role in modern machine learning, driving the field’s ever-expanding scaling of parameters, data and compute. While much research has gone into fitting scaling laws and predicting performance on pretraining losses and on discriminative evaluations such as multiple-choice question-answering benchmarks, comparatively little research has been done on fitting scaling laws and predicting performance on generative evaluations such as mathematical problem-solving or coding. In this work, we propose and evaluate three different pretraining scaling laws for fitting pass-at-$k$ on generative evaluations and for predicting pass-at-$k$ of the most expensive model using the performance of cheaper models. Our three scaling laws differ in the covariates used: (1) pretraining compute, (2) model parameters and pretraining tokens, (3) log likelihoods of gold reference solutions. We make four main contributions: First, we show how generative evaluations offer new hyperparameters (in our setting, $k$) that researchers can use to control the scaling laws parameters and the predictability of performance. Second, in terms of scaling law parameters, we find that the compute scaling law and parameters + tokens scaling law stabilize for the last $\mathord{\sim}1.5{-}2.5$ orders of magnitude, whereas the gold reference likelihood scaling law stabilizes for the last $\mathord{\sim}5$ orders of magnitude. Third, in terms of predictive performance, we find all three scaling laws perform comparably, although the compute scaling law predicts slightly worse for small $k$ and the log likelihoods of gold reference solutions predicts slightly worse for large $k$. Fourth, we establish a theoretical connection that the compute scaling law emerges as the compute-optimal envelope of the parameters-and-tokens scaling law. Our framework provides researchers and practitioners with insights and methodologies to forecast generative performance, accelerating progress toward models that can reason, solve, and create.
@@ 1 Introduction
Neural scaling laws, which predictably map resources to the performance of neural networks, are foundational for developing frontier AI systems (Kaplan et al., 2020b; Hoffmann et al., 2022b). Such empirical regularities for improving model performance from scaling parameters, data, and compute have driven pursuit of ever-larger models. Much of the research in neural scaling laws has focused on fitting and predicting model performance on pretraining losses, e.g., (Kaplan et al., 2020b; Hoffmann et al., 2022b; Clark et al., 2022; Hernandez et al., 2022; Sardana et al., 2024; Muennighoff et al., 2023; Porian et al., 2024; Pearce & Song, 2024) and on “downstream” discriminative evaluations like multiple-choice question-answering, e.g., (Schaeffer et al., 2023; Gadre et al., 2024; Schaeffer et al., 2024; Grattafiori et al., 2024; Chen et al., 2025; Bhagia et al., 2025).
While both pretraining losses and discriminative evaluations are undeniably useful, many capabilities we care about are generative: writing books, autoformalizing mathematics, or conducting cutting-edge research. Comparatively little research has investigated scaling laws for generative evaluations (OpenAI et al., 2024; Hu et al., 2024); for a deeper discussion of related work, please see Appendix A. Unlike pretraining losses and discriminative evaluations, generative evaluations introduce new considerations such as which metric is used to quantify performance and how samples are drawn from the model, including the sampling temperature (Ackley et al., 1985) and the sampling algorithm, e.g., top-k (Fan et al., 2018), top-p (Holtzman et al., 2020).
In this work, we propose a framework for fitting scaling laws and predicting performance on generative evaluations, focusing on a particular setting: benchmarks with verifiable binary rewards, with multiple attempts per problem, with performance scored using the “pass-at-kk” metric (Kulal et al., 2019; Chen et al., 2021). We chose this setting due to its widespread use in the literature (First et al., 2023; Brown et al., 2024; Hassid et al., 2024; Hughes et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Kwok et al., 2025). In this setting, we make the following contributions:
@@
    We reveal that a hyperparameter specific to this setting – the number of attempts per problem $k$ – offers a new lever to change the parameters of the scaling laws as well as how accurately the scaling laws extrapolate to more expensive models.
@@
    We quantify how stable each scaling law’s parameters are, finding that parameters for the compute scaling law and the parameters + tokens scaling law stabilize by the last $\mathord{\sim}1.5{-}2.5$ orders of magnitude of pretraining compute, whereas parameters for the gold reference likelihood scaling law are stable over the last $\mathord{\sim}5$ orders of magnitude.
@@
    We prove that the compute-only scaling law is the compute-optimal envelope of the parameters-and-tokens scaling law, obtained by minimizing the objective under a fixed compute budget. We also derive the optimal split of compute between model parameters and pretraining data, and a dimensionless misallocation penalty that explains when fixed training recipes underperform compute-optimal scaling.
Figure 1: Scaling of Pass Rates with Pretraining Compute (Full Fit). Each panel fits Eqn. 3 $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}$ to GSM8K pass rates for Pythia checkpoints across $\sim$5 orders of magnitude of pretraining compute. Panels show $k\in\{1,10^{2},10^{4}\}$. The fit curves capture the trend moderately well and reveal two effects of increasing $k$: (i) the irreducible error $\widehat{E}_{0}(k)$ falls from $2\to 0$ as $k$ increases from $1\to 100$; and (ii) the compute-dependent terms increase with $k$: the prefactor grows from $\widehat{C}_{0}(k=1)\approx 1e3$ to $\widehat{C}_{0}(k=10000)\!\sim\!7.8e5$ and the exponent rises from $\widehat{\alpha}(k=1)\!\approx\!0.120$ to $\widehat{\alpha}(k=100)\!\approx\!0.325$. Takeaway: Larger $k$ makes the relationship a pure power law and makes pass rates a steeper function of pretraining compute.
@@ 2 Methodology
Benchmark To study how language model performance changes during pretraining on generative evaluations, we used GSM8K (Cobbe et al., 2021), a well-established benchmark containing $8,500$ grade school math word problems designed to evaluate a model’s ability to perform multi-step mathematical reasoning. Each attempt at solving a problem yields a binary outcome: the attempt is either successful or unsuccessful. Due to limited budget, we used a randomly chosen subset of $128$ problems. See Appendix B for a description of how many samples we drew per problem per model.
Metric To evaluate the performance of each model, we used the ubiquitous pass-at-$k$ metric (Kulal et al., 2019). For the $i$-th problem, a model’s pass rate at $k$ attempts is the probability of the event that any of $k$ attempts are successful. Because generative evaluations sample from the language model, this per-problem pass rate must be estimated. We use the unbiased, low variance and numerically stable estimator introduced by Chen et al. (2021), which, for each problem, draws $n_{i}>k$ samples, counts the number of successes $s_{i}\in[0,n_{i}]$, and then averages over all subsets of size $k$:
$\mathrm{pass_{i}@k}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad 1-\dbinom{n_{i}-s_{i}}{k}/\dbinom{n_{i}}{k}$. (1)
@@
Pass rates are sometimes phrased as the fraction of problems in the benchmark solved (“coverage”), but these two quantities are mathematically equivalent (Brown et al., 2024; Schaeffer et al., 2025b).
Figure 2: Scaling Law Parameters Vary with Attempts per Problem (Full Fit). In comparison with pretraining loss scaling laws and discriminative evaluation scaling laws, the number of attempts per problem $k$ introduces a new hyperparameter that affects the parameters of generative evaluation scaling laws, and also how predictable $\mathrm{pass_{\mathcal{B}}}@k$ is. Here we show the effect of $k$ for the pretraining compute scaling law (Eqn. 3); for a side-by-side comparison with our next two scaling laws, see Appendix C. Left: The irreducible error $E_{0}(k)$ falls roughly exponentially with $k$ and is $\approx 0$ by $k\!\approx\!10^{2}$, consistent with the probability that a problem remains unsolved decaying exponentially in the number of attempts. Center: The compute prefactor $C_{0}(k)$ increases monotonically with $k$, indicating that once $E_{0}(k)\!\to\!0$, variation is dominated by the compute-dependent term. Right: The compute exponent $\alpha(k)$ increases smoothly, from $\sim\!0.15$ at k=1k{=}1 to $\sim\!0.3$ at $k{=}10^{4}$, implying steeper and more regular scaling at larger $k$. Hue corresponds to $k$.
Sampling We used temperature-only sampling (Ackley et al., 1985) at $\tau{=}1.0$ and did not sweep on account of prior work (Schaeffer et al., 2025a).
Language Model Family For our experiments, we used the Pythia family (Biderman et al., 2023), which contains 8 models from 14M to 12B parameters pretrained on 300B tokens from The Pile (Gao et al., 2020). We used 8 checkpoints per parameter size of the non-deduplicated variants, but excluded extremely overtrained checkpoints identified as abnormal by prior work (Godey et al., 2024). Each model checkpoint has a specific number of model parameters ($N$) and has been trained on a certain number of tokens ($D$). We took the standard approximation for pretraining compute of $C\approx 6\cdot N\cdot D$ (Kaplan et al., 2020b; Hoffmann et al., 2022b; Pearce & Song, 2024; Gadre et al., 2024; Schaeffer et al., 2024; Porian et al., 2024). We use intermediate checkpoints in lieu of fully trained models, and note that our results would likely improve with greater experimental control.
@@ 3.1 Fitting Pass Rates from Pretraining Compute
Motivated by the GPT-4 Technical Report (OpenAI et al., 2024), we posited that the negative log of benchmark pass rates follows a scaling law as a function of model pretraining compute $C$:
@@
Figure 3: Predicting Pass Rates from Scaling Pretraining Compute (Backtesting). We evaluate predictability via backtesting: we iteratively fit Eq. 3 on subsets of models up to a maximum pretraining compute, then extrapolate to predict the most expensive model’s $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)$. The most expensive model is Pythia 12B-parameter 300B-token ($\approx 2.16\times 10^{22}$ FLOP). The $x$-axis in all panels is the ratio of the largest-compute checkpoint included in the fit to the target model’s compute. Top Left: Relative error decreases as higher-compute checkpoints are included and then plateaus: for $k\!\in\!\{1,10^{2}\}$ the plateau occurs once the fit includes checkpoints within $\mathord{\sim}$2 orders of magnitude of the target; for $k\!=\!10^{4}$ it occurs within $\mathord{\sim}1.5$ order. Other Three Panels: Backtested estimates of the scaling law parameters converge as the fit includes higher-compute checkpoints, stabilizing to full-fit values once included models are within $\mathord{\sim}2$ orders of magnitude of the target.
We fit the scaling law parameters based on the model checkpoints’ pretraining compute and their corresponding $\mathrm{pass_{\mathcal{B}}}@k$. Visually, the fit scaling laws captured the trend reasonably well for different values of number of attempts per problem $k$ (Fig. 1). We next evaluated what role the number of attempts per problem $k$ has on the scaling law parameters: As $k$ increases, the irreducible error term $E_{0}(k)$ falls roughly exponentially with $k$ and is effectively 0 by $k\approx$1\text{\times}{10}^{2}$\$ (Fig. 2 Left). This is consistent with the probability a problem goes unsolved falling exponentially with the number of attempts per problem (Levi, 2025; Schaeffer et al., 2025b). Thus, by increasing $k$, Eqn. 3 becomes a pure power law with no irreducible error. In comparison, as $k$ increases, the compute prefactor and compute exponent also increase, with the compute prefactor rising $\mathord{\sim}4$ orders of magnitude (Fig. 2 Center) and the compute exponent rising moderately from $1.20\text{\times}{10}^{-1}$ to $3.25\text{\times}{10}^{-1}$ (Fig. 2 Right).
@@ 3.2 Predicting Pass Rates from Pretraining Compute
Figure 4: Scaling of Pass Rates with Model Parameters and Pretraining Data (Full Fit). Each panel fits Eqn. 4, $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(N,D,k)=\mathcal{E}_{0}(k)+N_{0}(k)\,N^{-\beta(k)}+D_{0}(k)\,D^{-\gamma(k)}$. Relative to the compute scaling law, using parameters $N$ and tokens $D$ instead yields tighter in-range fits for all $k$. Consistent with Fig. 2, the irreducible error decreases sharply with $k$ (e.g., ^0: $\widehat{\mathcal{E}}_{0}\!:\ 3.87\!\to\!$0 by $k\!\approx\!10^{2}$), after which variation is dominated by the $(N,D)$ terms. Despite the better global fit, the largest-compute model checkpoint in each panel exhibits comparatively large relative error.
@@ 4.1 Fitting Pass Rates from Model Parameters and Pretraining Tokens
Figure 5: Predicting Pass Rates from Scaling Parameters and Tokens (Backtesting). We evaluated how accurately the parameters + tokens scaling law (Eqn. 4) predicts the most expensive model’s $\log\!(\mathrm{pass}_{\mathcal{B}}@k)$. Top Left: The relative error decreases as higher-compute checkpoints are included, plateauing once the fit includes models within $\mathrm{\sim}2$ orders of magnitude of the target. Other Five Panels: The estimates of the five scaling law parameters are initially unstable but converge to their full-fit values once included models are within $\mathord{\sim}2$ orders of magnitude of the target.
We fit this parameters + tokens scaling law. Visually, this law provides a better characterization of performance for the full range of models, with the fitted curves appearing tighter and reducing the residual scatter (Fig.4), compared against the pretraining compute law (Fig.1). Consistent with our findings for the compute-only law, the irreducible error term $\widehat{\mathcal{E}}_{0}(k)$ decreases with the number of attempts per problem $k$, although it does so more gradually, reaching zero by $k\approx$3\text{\times}{10}^{2}$\$ (Fig. 10). However, despite the improved global fit, we observed that the largest-compute model checkpoints in our dataset exhibit a comparatively large relative error, suggesting that while this law better explains the performance of less-trained models, it may struggle with extrapolation to the frontier.

### v2

@@ Abstract
Neural scaling laws have driven the field’s ever-expanding exponential growth in parameters, data and compute. While scaling behaviors for pretraining losses and discriminative benchmarks are well established, generative benchmarks such as mathematical problem-solving or software engineering remain under-explored. We propose and evaluate three different pretraining scaling laws for fitting pass-at-$k$ on generative evaluations and for predicting pass-at-$k$ of the most expensive model using cheaper models. Our three scaling laws differ in the covariates used: (1) pretraining compute, (2) model parameters and pretraining tokens, (3) log likelihoods of gold reference solutions. First, we demonstrate that generative evaluations introduce new hyperparameters (in our setting, $k$) that act as a control lever for scaling behavior, modulating both the scaling law parameters and the predictability of performance. Second, we identify a stark difference in parameter stability: while the compute and parameters+tokens laws stabilize for only the last $\mathord{\sim}1.5\mathord{-}2.5$ orders of magnitude, the gold reference likelihood law is uniquely stable, converging across $\mathord{\sim}5$ orders. Third, in terms of predictive performance, we find all three scaling laws perform comparably, although the compute law predicts slightly worse for small $k$ and the gold reference law predicts slightly worse for large $k$. Finally, we establish a theoretical connection, proving that the compute scaling law emerges as the compute-optimal envelope of the parameters-and-tokens law. Our framework provides researchers and practitioners with insights and methodologies to forecast generative performance, accelerating progress toward models that can reason, solve, and create.
@@ 1 Introduction
Neural scaling laws, which predictably map resources to the performance of neural networks, are foundational for developing frontier AI systems (Kaplan et al., 2020b; Hoffmann et al., 2022b). Such empirical regularities for improving model performance from scaling parameters, data, and compute have driven pursuit of ever-larger models. Much of the research in neural scaling laws has focused on fitting and predicting model performance on pretraining losses, e.g., (Kaplan et al., 2020b; Hoffmann et al., 2022b; Clark et al., 2022; Hernandez et al., 2022; Sardana et al., 2024; Muennighoff et al., 2023; Porian et al., 2024; Pearce and Song, 2024) and on “downstream” discriminative evaluations like multiple-choice question-answering, e.g., (Schaeffer et al., 2023; Gadre et al., 2024; Schaeffer et al., 2024b; Grattafiori et al., 2024; Chen et al., 2025; Bhagia et al., 2025).
While pretraining losses and discriminative evaluations are undeniably useful, many capabilities we care about are generative: writing books, autoformalizing mathematics, or conducting cutting-edge research. Scaling laws for generative tasks remain undercharacterized (Appendix A) (but see (OpenAI et al., 2024a; Hu et al., 2024)); for example, Gadre et al. (2024) studied 46 tasks, none of which were generative. Generative evaluations differ from discriminative evaluations in many ways, most crucially that performance is calculated from the model’s open-ended samples, which introduces new considerations such as how attempts are sampled from the model, including the sampling temperature and the sampling algorithm, and which metric is used to measure performance.
In this work, we propose a framework for fitting scaling laws and predicting performance on generative evaluations, focusing on a particular setting: benchmarks with verifiable binary rewards, with multiple attempts per problem, with performance scored using the “pass-at-kk” metric (Kulal et al., 2019; Chen et al., 2021). We chose this setting due to its widespread use in the literature, e.g., (First et al., 2023; Brown et al., 2024; Hassid et al., 2024; Hughes et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Kwok et al., 2025). In this setting, we make the following contributions:
@@
    We reveal that a hyperparameter specific to this setting – the number of attempts per problem $k$ – offers a new lever to shape the scaling and predictability of model performance.
@@
    We quantify the stability of each scaling law’s parameters. We find that parameters for the compute scaling law and the parameters + tokens scaling law stabilize by the last $\mathord{\sim}1.5{-}2.5$ orders of magnitude of pretraining compute, whereas parameters for the gold reference likelihood scaling law are stable over the last $\mathord{\sim}5$ orders of magnitude.
@@
    We prove that the compute-only scaling law is the compute-optimal envelope of the parameters-and-tokens scaling law, obtained by minimizing the objective under a fixed compute budget. From this, we derive a dimensionless misallocation penalty that explains quantitatively when and why pretraining recipes underperform compute-optimal scaling.
Figure 1: Scaling of GSM8K Pass Rates with Pretraining Compute (Full Fit). Each panel fits Eqn. 3 $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}$ to GSM8K pass rates for Pythia checkpoints across $\sim$5 orders of magnitude of pretraining compute. Increasing $k$ drives two major shifts: (i) The irreducible error $E_{0}(k)$ vanishes (from $\approx 2\to 0)$, removing the performance plateau, and (ii) the power law steepens, with the exponent $\alpha(k)$ rising from $\approx\!0.121\to\!0.375$. Takeaway: Larger $k$ eliminate irreducible error and yield steeper pass-at-$k$ rates with respect to pretraining compute.
@@ 2 Methodology
Language Model Family: For our experiments, we used the Pythia family (Biderman et al., 2023) of 8 models from 14M to 12B parameters pretrained on up to 300B tokens from The Pile (Gao et al., 2020). We restricted our analysis to Pythia family because it is, to our knowledge, the only public model family meeting three criteria: (1) densely sampled across both model parameters ($N$) and pretraining tokens ($D$); (2) spanning many orders of magnitude in pretraining compute (5); and (3) known token budgets per checkpoint. Other families (e.g., Llama 3, Qwen 3) either lack sufficient intermediate checkpoints to robustly fit a 5-parameter scaling law or do not disclose specific per-model pretraining token budgets. We used 8 checkpoints per parameter size of the non-deduplicated variants, but excluded extremely overtrained checkpoints identified as abnormal (Godey et al., 2024). We approximated pretraining compute as $C\approx 6\,N\,D$ (Kaplan et al., 2020b; Hoffmann et al., 2022b; Pearce and Song, 2024; Gadre et al., 2024; Schaeffer et al., 2024b; Porian et al., 2024) and used intermediate checkpoints in lieu of fully trained models.
Benchmarks: We used two well-established generative tasks: GSM8K (Cobbe et al., 2021) with $8,500$ grade school math word problems, and MATH (Hendrycks et al., 2021) with $12,500$ competition mathematics problems. Each attempt at solving a problem yields a binary outcome: success or failure. Because generative evaluations are computationally expensive, to enable a rigorous analysis across 5 orders of magnitude of compute with high sample counts, we evaluated all checkpoints on a randomly selected subset of 128 problems from GSM8K and 128 problems from MATH; see Appendix B for the number of samples per problem per model, and Appendix C for the costs.
Metric: We measured the performance of each model using the pass-at-$k$ metric (Kulal et al., 2019), defined as the probability that if a model makes $k$ attempts on the $i$th problem, at least one attempt is successful. We estimate this per-problem pass rate using the unbiased, low variance and numerically stable estimator of Chen et al. (2021), which, for each problem, draws $n_{i}>k$ samples, counts the number of successes $s_{i}\in[0,n_{i}]$, and then averages over all subsets of size $k$:
$\mathrm{pass_{i}@k}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad 1-\frac{\tbinom{n_{i}-s_{i}}{k}}{\tbinom{n_{i}}{k}}$ (1)
@@
Pass rates are sometimes phrased as the fraction of problems in the benchmark solved (“coverage”), which is mathematically equivalent (Brown et al., 2024; Schaeffer et al., 2025b). Importantly, pass-at-$k$ is fundamentally different from (discriminative) accuracy, as pass-at-$k$ is a continuous probability derived from the model’s generative distribution whereas accuracy is a discriminative function of the correct and incorrect multiple choices (Appendix A).
Figure 2: Number of Attempts per Problem $k$ Shapes Scaling Law Parameters (Full Fit). We fit Eqn. 3 to each $k$ (hue); see Appendix D for the next two scaling laws. Left: Irreducible error $E_{0}(k)$ decays roughly exponentially with $k$ and is $\approx 0$ by $k\!\approx\!10^{2}$. Center: Compute prefactor $C_{0}(k)$ increases monotonically with $k$, indicating that once $E_{0}(k)\!\to\!0$, the compute-dependent term dominates. Right: Compute exponent $\alpha(k)$ increases smoothly, from $\sim\!0.15$ at k=1k{=}1 to $\sim\!0.3$ at $k{=}10^{4}$, indicating that larger sampling budgets yield steeper, more favorable scaling behaviors.
Generation: Generative evaluation introduces a vast hyperparameter space (e.g., temperature (Ackley et al., 1985), top-p (Holtzman et al., 2020), top-k (Fan et al., 2018)). We deliberately isolated the number of attempts per problem ($k$) as our primary variable because it is a simple yet effective lever for scaling inference-time compute (OpenAI et al., 2024b; Snell et al., 2024; Brown et al., 2024; DeepMind et al., 2025; DeepSeek-AI et al., 2025a) and because previous work has studied the effect of sampling algorithm and sampling temperature on GSM8K (Chen et al., 2021; Schaeffer et al., 2025a). We used temperature-only sampling at $\tau=1.0$, holding the decoding algorithm fixed.
Figure 3: Predicting GSM8K Pass Rates from Scaling Pretraining Compute (Backtesting). We evaluate predictability via backtesting: iteratively fitting Eq. 3 on subsets of models $(C\leq C_{\mathrm{max}})$ to predict the most expensive model (Pythia 12B-parameter 300B-token; $\approx 2.16\times 10^{22}$ FLOP)’s $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)$. The $x$-axis denotes the compute horizon relative to the target $(C_{\mathrm{max}}/C_{\mathrm{target}})$. Top Left: Relative error decreases and then plateaus: reliable prediction requires checkpoints within $\mathord{\sim}$2 orders of magnitude of the target for $k\in\{1,10^{2}\}$ and $\mathord{\sim}1.5$ for $k=10^{4}$. Other Three Panels: Backtested estimates of the scaling law parameters are initially unstable but converge to their full-fit values once fits include models within $\mathord{\sim}2$ orders of magnitude of the target.
@@ 3.1 Fitting Pass Rates from Pretraining Compute
Motivated by the GPT-4 Technical Report (OpenAI et al., 2024a), we posited that the negative log of benchmark pass rates follows a scaling law as a function of model pretraining compute $C$:
@@
We fit the scaling law parameters based on the model checkpoints’ pretraining compute and their corresponding $\mathrm{pass_{\mathcal{B}}}@k$. Visually, the fit scaling laws captured the trend reasonably well for different values of number of attempts per problem $k$ (Fig. 1). We next evaluated what role the number of attempts per problem $k$ has on the scaling law parameters: As $k$ increases, the irreducible error term $E_{0}(k)$ falls roughly exponentially with $k$ and is effectively 0 by $k\approx$1\text{\times}{10}^{2}$\$ (Fig. 2 Left). This is consistent with the probability a problem goes unsolved falling exponentially with the number of attempts per problem (Levi, 2025; Schaeffer et al., 2025b). Thus, by increasing $k$, Eqn. 3 becomes a pure power law with no irreducible error. In comparison, as $k$ increases, the compute prefactor and compute exponent also increase, with the compute prefactor rising $\mathord{\sim}4$ orders of magnitude (Fig. 2 Center) and the compute exponent rising moderately from $1.21\text{\times}{10}^{-1}$ to $3.75\text{\times}{10}^{-1}$ (Fig. 2 Right).
@@ 3.2 Predicting Pass Rates from Pretraining Compute
Figure 4: Scaling of GSM8K Pass Rates with Parameters and Tokens (Full Fit). Each panel fits Eqn. 4, $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(N,D,k)=\mathcal{E}_{0}(k)+N_{0}(k)\,N^{-\beta(k)}+D_{0}(k)\,D^{-\gamma(k)}$. Decomposing compute into parameters $N$ and tokens $D$ instead yields tighter in-range fits than the compute law for all $k$. Consistent with Fig. 2, the irreducible error decreases sharply with $k$ (0: $\mathcal{E}_{0}\!:\ 3.87\!\to\!$0 by $k\!\approx\!300$), after which variation is dominated by $(N,D)$ terms. However, despite the better global fit, the largest-compute model checkpoint in each panel exhibits comparatively large relative error.
@@ 4.1 Fitting Pass Rates from Model Parameters and Pretraining Tokens
Figure 5: Predicting GSM8K Pass Rates from Scaling Parameters and Tokens (Backtesting). We evaluate how accurately the parameters + tokens scaling law (Eqn. 4) predicts the most expensive model’s $\log(\mathrm{pass}_{\mathcal{B}}@k)$. Top Left: The relative error decreases as higher-compute checkpoints are included, plateauing once the fit includes models within $\mathrm{\sim}2$ orders of magnitude of the target. Other Five Panels: Estimates of the five scaling law parameters are initially unstable but converge to their full-fit values once included models are within $\mathord{\sim}2$ orders of magnitude of the target.
We fit this parameters + tokens scaling law. Visually, this law provides a better characterization of performance for the full range of models, with the fitted curves appearing tighter and reducing the residual scatter (Fig.4), compared against the pretraining compute law (Fig.1). Consistent with our findings for the compute-only law, the irreducible error term $\widehat{\mathcal{E}}_{0}(k)$ decreases with the number of attempts per problem $k$, although it does so more gradually, reaching zero by $k\approx$3\text{\times}{10}^{2}$\$ (Fig. 12). However, despite the improved global fit, we observed that the largest-compute model checkpoints exhibit comparatively large relative error, suggesting that while this law better explains the performance of cheaper models, it may struggle with extrapolation to the frontier.